OpenStudy (narissa):

which is an equation in point-slope form for the given point and slope. Point: (8,-4 ); Slope: -5/6

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (narissa):

same thing okay i got y-4=5/6(x-8)

pooja195 (pooja195):

two negatives next to each other make a positive try again :)

OpenStudy (narissa):

y-4=-5/6(x+8)?

pooja195 (pooja195):

I got lagged out of the site sorry No it's still incorrect y-y1=m(x-x1) y1=? x1=?

OpenStudy (skullpatrol):